OpenStudy (anonymous):

if q is inversly proportional to the square of r and q is 4/81 when r is 9, find q when r is 2.

OpenStudy (rc94):

Firstly take q=a/r^2 as a general form

OpenStudy (rc94):

Now substitute in q=4/81 and r=9 so you have the equation: 4/81 = a/9^2 Ie 4/81 = a/81 and solve for a. Then use this value of a for the last part

OpenStudy (anonymous):

k